OpenStudy (anonymous):

is 2x-5y==7 linear or not

OpenStudy (anonymous):

2x-5y=7?

OpenStudy (anonymous):

you can use algebra to solve for y - notice that none of the variables have a higher power than 1

OpenStudy (anonymous):

or a power other than 1

OpenStudy (anonymous):

a linear equation always take the form y= mx+b

OpenStudy (anonymous):

good luck :)

OpenStudy (amistre64):

i linear equation satisfies the following: f(a+b) = f(a) + f(b) f(ca) = c f(a) for some constant c

OpenStudy (amistre64):

y = mx; is a linear equation. y = mx + b; b not= 0, is an affine equation